Required information The following information applies to the questions displayed below. Ramirez Company installs a computerized manufacturing machine in its factory at the beginning of the year at a cost of $45,300. The machines useful life is estimated at 10 years, or 403.000 units of product, with a $5,000 salvage value. During its second year, the machine produces 34,300 units of product Determine the machines second-year depreciation using the double-declining-balance method. Double-declining-balance Depreciation Annual Depreciation Choose Factors: Choose Factor(%) Depreciation expense First years depreciation Second years depreciation
Answer #1
Double declining depreciation method compute depreciation based on 100%/ number of useful life *2 as depreciation rate.
Salvage value is not considered under double declining depreciation method.
Cost of machine 45300
use full life 10 year
Depreciation rate = =100%/10*2 20%
Choose factors * Choose factors (%) = Annual depreciation
First year depreciation 45300 20% 9060.00
Second year depreciation 36240 20% 7248.00
First year depreciation = 45300*20% = 9060
Second year depreciation = (45300-9060)*20% = 7248